OpenStudy (shreenithaa):

express each conjunction as a single mathematical statement. 5 is less than or equal to 3a-4 and 3a-4<17

OpenStudy (shreenithaa):

is it 5 is less than or equal to 3a-4<17 +4+4+4 9 is less than or equal to 3a <21 ////////////////////////// 3 3 is less than or equal to a<7

OpenStudy (shreenithaa):

is this right?

zepdrix (zepdrix):

I don't think they wanted you to do that far with it. You're told that \(\large\rm 5\le~stuff\) and \(\large\rm stuff\lt 17\) Therefore, \(\large\rm 5\le~stuff\lt17\).

zepdrix (zepdrix):

I think they wanted you to stop there, not solve for a :o

OpenStudy (shreenithaa):

s but my teacher likes things simplified

OpenStudy (shreenithaa):

so i am right?

zepdrix (zepdrix):

Oh you got 3 and 7? :) Ah good, those looks correct! \(\large\rm 3\le a\lt 7\)

OpenStudy (shreenithaa):

how do u do that less than or equal to and greater than or equal to sign

OpenStudy (jb1625):

zepdrix can you help me with this? Points R, S, and T are plotted on a number line. The coordinate of point R is –6. Points S and T are the same distance from point R. What could be the coordinates of S and T? A. S = 0, T = 6 B. S = –4, T = –2 C. S = –9, T = –12 D. S = –10, T = –2

zepdrix (zepdrix):

@JB1625 please don't hijack someone else's question. :o Go open your own thread.

zepdrix (zepdrix):

In text you would write it like this `<=`, In the equation tool, you have to type \le or you can find it on the list of symbols.

OpenStudy (jb1625):

Well I just wanted to make sure that you saw it sorry

OpenStudy (shreenithaa):

ooh thnx and its k

zepdrix (zepdrix):

Use the pager program @JB1625 :) If you type @zepdrix somewhere in your own thread, it will alert me to it.

zepdrix (zepdrix):

Or you could google `less than or equal to` and copy paste the symbol :) That might be easier. 3 ≤ a < 7

OpenStudy (shreenithaa):

thnx i was confused at first how others did it now i got it